# positioning

### v1

@@ Abstract
Rather than having each newly deployed robot create its own map of its surroundings, the growing availability of SLAM-enabled devices provides the option of simply localizing in a map of another robot or device. In cases such as multi-robot or human-robot collaboration, localizing all agents in the same map is even necessary. However, localizing e.g. a ground robot in the map of a drone or head-mounted MR headset presents unique challenges due to viewpoint changes. This work investigates how active visual localization can be used to overcome such challenges of viewpoint changes. Specifically, we focus on the problem of selecting the optimal viewpoint at a given location. We compare existing approaches in the literature with additional proposed baselines and propose a novel data-driven approach. The result demonstrates the superior performance of the data-driven approach when compared to existing methods, both in controlled simulation experiments and real-world deployment.
@@ I Introduction
As an example of this larger question, this paper delves into the specific scenario of localizing a ground robot in apre-existing map generated from a mixed-reality headset. This application holds considerable practical value, as it obviates the need to re-map an entire building if a suitable map already exists. Furthermore, for seamless collaboration between human-robot teams, the ability to localize mixed-reality devices and robots within a shared map becomes imperative, as highlighted in previous studies [6, 10, 25]. However, accurately localizing a robot in a map created with a head-mounted camera rig introduces its own challenges. These challenges stem from the use of diverse sensor devices and are compounded by significant variations in viewpoint between the mapping trajectory and the operational height of the robot. Such variations result in a reduced visual overlap between the robot’s perspective and the map, subsequently diminishing the localization performance. This issue becomes particularly prominent in the case of ground robots, including quadrupeds, where obstacles such as chairs, tables, and furniture frequently obscure substantial portions of the robot’s environment.
@@
[…]
Figure 1: Selected Viewing Directions of three methods to visually localize at the same given location with respect to the landmarks in red. The passive strategy of looking forwards (top left) results in a localization error of 3.01m / 106deg. A strategy inspired by [8] to maximize the similarity with the viewing angle towards the landmark during mapping localizes with 0.15m / 0.8deg error. Finally, our data-driven viewpoint transformer (VPT) (top right) achieves 0.01m / 0.1deg localization accuracy.
In contrast to mixed-reality devices, robots possess the valuable ability to autonomously select their own viewpoints. This leads us to investigate whether active viewpoint selection can effectively address the challenges associated with cross-agent visual localization. In literature, this concept is widely recognized as active perception, and within our specific context, it is referred to as Active Visual Localization. The core objective of Active Visual Localization is to determine the pose of a camera within an existing map representation of the environment. One common approach involves assessing the localization utility of various viewpoints, typically through metrics like the Fisher Information metric or a combination of hand-crafted heuristics.. While extensive work has have been directed towards integrating these utility calculations into planning frameworks, a relatively less studied component is how the utility value itself can be calculated, particularly in scenarios where the unique challenges discussed earlier.
This work delves into the exploration of a effective utility function to actively enhance visual localization of a robot in a within that was created from a different perspective. Our primary focus lies in evaluating established viewpoint selection and assessment criteria while introducing a novel data-driven approach to viewpoint scoring. The key contributions of this paper are:
@@
  • a comparison and thorough evaluation of viewpoint selection methods for the problem of visual localization between heterogenous agents
  • a real-world validation of our findings by integrating the viewpoint selection into an active viewpoint planner for a robot with an arm-mounted camera
We further open-source our implementations and collected datasets.
Figure 2: Difference in perspective between a head-mounted sensor rig used for mapping (left) and a ground robot (right).
@@ II Related Work
Active Vision describes the case where an agent has the ability to move visual sensors with the goal of improving the performance of perceptual tasks [2]. This concept finds application across various domains where visual information is utilized, such as exploration [30], data collection [27], or active learning [38]. The case studied in this work focuses on the task of improving visual localization by choosing the most informative viewpoint at a given position, which can be considered a subset of Active Vision: Active Visual Localization. Many existing works in the literature study how informative metrics can be incorporated into motion planning frameworks with the goal of optimizing a motion to maximize localization accuracy [26, 7, 35]. Within the scope of this research, our focus is on the specific task of augmenting visual localization by selecting the most informative viewpoint at a given position.
@@
Regarding viewpoint selection in the realm of active vision,early approaches primarily relies on handcrafted metrics to gauge the uncertainty or reliability of a given viewpoint in contributing to the system’s state  [28, 34, 15, 13, 36, 23]. Some methods take a different path by deriving evaluation metrics from the metric map and constructing a global utility map.. Authors of  [37] propose a solution rooted in Fisher information theory  [12, 20]. The central task they explore consists of determining the amount of information a viewpoint from a given pose will contribute to the localization process. They develop a novel map representation that enables efficient computation of the Fisher information for 6-DoF visual localization, known as the Fisher Information Field. . Similar ideas of using Fisher Information for viewpoint selection have also been explored in recent works[17, 16, 1]. However, it’s important to note that these metrics often rely on heuristics, necessitate the design of specific handcrafted utility functions, and may have limitations in their representation capabilities for diverse and complex scenarios.
@@
Another category of works introduces additional vision tasks to enhance the metric extraction process, notably incorporating semantic information [33]. For instance, the work of [3] aims to improve navigation performance by including semantic information, in order to discern perceptually informative areas of the environment. Such kind of work enriches the semantic understanding capability, however highly relies on the performance of the semantic module, and usually requires prior knowledge to link certain semantic classes that clearly correlate to visual informativeness.
@@
A separate line of research focus on using data-driven approaches to active visual localization. These works mostly choose to formulate the problem as a reinforcement learning task, tightly integrating metric determination with robot execution [4, 5, 11]. As an example, [18] trains an information-aware policy to find traversable paths as well as reduce the uncertainty of the environment. The learning-based model significantly enhance the robot’s comprehension of its surroundings. However, it’s worth noting that these models typically demand substantial computational resources for training and may require the simplification of the environment model to prevent overfitting.
@@ III Method
Problem Statement: Let $\bm{x}=(\bm{p},\bm{q})$, where $\bm{p}\in\mathcal{R}^{3}$ and $\bm{q}\in so(3)$ are the position and orientation of the robot sensor. Given a map representation of the environment $\mathcal{M}$ and an initial guess of the robot pose $\hat{\bm{x}}$, the goal of the active viewpoint planning is to find an orientation for the robot sensor, at a certain position, such that the visual localization method returns the most accurate location estimation $\bar{\bm{x}}$, i.e.,
@@
[…]
In this paper, we choose $\mathcal{M}$ combines the landmark point cloud $\mathcal{M}_{l}$ and the TSDF $\mathcal{M}_{t}$. Any viewpoint selection policy $\uppi(\cdot)$ then takes $\mathcal{M}=(\mathcal{M}_{l},\mathcal{M}_{t})$ as prior knowledge of the environment, and selects a viewpoint for a certain position:
$\bm{q}$ $=\uppi(\mathcal{M},\hat{\bm{p}})$ (3)
For example, we implement the idea of Fisher Information policy of [37] as:
@@
where $v(\cdot)$ is the binary visibility of landmark $i$ and $\textrm{I}_{i}$ is the Fisher Information Metric of observing landmark $i$. To determine the visibility $v(\cdot)$, we use the TSDF map $\mathcal{M}_{t}$. For each sampled viewpoint, a raycasting step is performed, during which we filter out the landmarks that are in collision with the environment, with a small tolerance to account for inaccuracies in the process.
@@
where $\uppi_{\textrm{max}}$ selects simply the view with the maximum visible landmarks, and $\uppi_{\textrm{angle}}$ uses a stricter visibility criterion inspired by [8] that only considers a landmark visible if its relative location to $\bm{x}$ is similar to those poses from which it was seen during mapping. This is illustrated in Figure 3.
Figure 3: Illustration of the Angle criteria. Given poses that see the landmark (red dot) during mapping, the green pose is within the directional range and has a higher chance of re-detecting the landmark. The red pose however sees the landmark from a new direction and does not count it based on the criteria.
Data-driven viewpoint scoring On top of these baselines, we propose a data-driven approach. In this approach, we gather the following essential information for each viewpoint:
  • distance to every landmark
  • minimum and maximum distance per landmark during mapping
  • viewing angle for every landmark
  • minimum and maximum viewing angle per landmark during mapping
@@
  • number of landmarks in the previously seen angle range
  • corresponding DINO [22] appearance features for every landmark in the current viewpoint
We build a data collection pipeline in simulated environments. To collect training data, we record for every viewpoint the above information, as well as the ground truth pose and the result of the visual localization method at that viewpoint. Based on this data, we train neural networks to classify whether a viewpoint has been localized within an error threshold. In particular, we distinguish between a MLP-based method $\uppi_{MLP}$ and a transformer-based method $\uppi_{VPT}$. The difference between these is the input encoding of the collected information. We design lightweight models in order to maintain real-time capability. Both models end with a softmax layer to determine the viewpoint score.
Because an MLP requires a fixed input dimensionality, but the number of landmarks varies across different viewpoints, we set a feature aggregation step in $\uppi_{MLP}$. For each information in the list above, we build a histogram to aggregate the corresponding value of all the filtered landmarks. For the pixel coordinates, we aggregate the information in a 2D heatmap instead of a 1D histogram. DINO features cannot be processed in the MLP.
Our main motivation to investigate a transformer architecture is its ability to process input data of variable length. Thus, inputs can be provided in a per-landmark fashion, which also allows for the inclusion of features that cannot be easily aggregated, such as DINO appearance per landmark.
Once the networks are trained, for a given $\bm{p}$, we evaluate all candidate viewpoints, assigning a localization score for each candidate, and then choosing the one with the highest score, e.g.:
@@
where $\mathbf{F}$ are the input features, and $f_{VPT}(\cdot)$ returns the output score of the transformer-based classifier.
@@ IV Experiments
We build a complete data collection pipeline for training, validation, and testing in both simulated and real-world, and conduct a series of experiments. We also implement several baseline methods for comparison. The data collection and model training is mostly done within the simulated world, details can be found in IV-A and 4. All the baseline methods and our evaluation strategy are introduced in IV-C. In the end, we show our real-world experiment in IV-D
@@ IV-A Data Generation
Simulated Scenario In this work, we focus on indoor scenarios, and we choose a selection of scenes from the Habitat-Matterport 3D (HM3D) dataset [24]. HM3D provides high-quality 3D reconstructions of real-world indoor environments with textured 3D meshes. We choose 9 scenes, shown in Figure 4, that we import into a simulator. In the simulator, we then manually record controlled trajectories with a simulated camera rig resembling the Microsoft HoloLens 2 [14] at a ”human-like” height.
Data Collection To obtain $\mathcal{M}_{t}$, we could use the depth image of the sensors and perform TSDF integration. However, Isaac Sim provides a simpler way to directly obtain occupancy information of a 3D scene using its built-in function [^1] https://docs.omniverse.nvidia.com/isaacsim/latest/advanced_tutorials/
tutorial_advanced_occupancy_map.html
, which generates accurate occupancy maps for us to perform conclusion and collision checking.
To obtain $\mathcal{M}_{l}$, we use the hloc/COLMAP [29, 31] mapping and localization framework to extract 2D local features from the image and triangulate them into 3D landmarks. We also use hloc as our $\textit{loc}(\cdot)$ in eq. 2. The collected images from the simulated HoloLens are then fed into the hloc pipeline to build the landmarks map. The per-landmark features in the feature vector $\mathbf{F}$ are also created and attached to each landmark at this stage. Additionally, landmarks that do not fit the following criteria are removed with the aim of improving map quality:
  • observed in at least 4 mapping images
  • observation distance within the range [0.1m, 8.0m]
To recover the scale of the map, one option would be to create the reconstruction with known camera poses. However, this would not be representative of how this process could be done using real hardware, where poses would be obtained from some form of VI-SLAM module. Thus, in our experiment, the maps are created without the exact camera poses, instead building the reconstructions from the images alone and then using RANSAC to estimate the scale with respect to the simulator poses.
To collect training data, we generate random camera trajectories at a ”robot-like” height with a simulated single RGB-D camera, which serves as an onboard sensor of the robot. At each waypoint, we take a group of viewpoint samples, and save its image observations, together with the corresponding ground-truth pose. As shown in Equations 1 and 2, we then run hloc for each viewpoint sample using its image observation and calculate the difference between the estimation pose and the ground truth.
@@
TABLE I: Evaluation of viewpoint strategies in the testing environments using both SuperPoint(Left) and SIFT(Right) local features. Recall percentages are shown at varying distance and orientation thresholds, highlighting the best and second-best performing methods. Rows marked as ‘with occlusion filter’ are those where landmarks are filtered with considering the occlusion. The bottom row is an orcale method that selects the smallest possible error among sampled viewpoints.

### v2

# Active Visual Localization for Multi-Agent Collaboration: A Data-Driven Approach

@@ Abstract
Rather than having each newly deployed robot create its own map of its surroundings, the growing availability of SLAM-enabled devices provides the option of simply localizing in a map of another robot or device. In cases such as multi-robot or human-robot collaboration, localizing all agents in the same map is even necessary. However, localizing e.g. a ground robot in the map of a drone or head-mounted MR headset presents unique challenges due to viewpoint changes. This work investigates how active visual localization can be used to overcome such challenges of viewpoint changes. Specifically, we focus on the problem of selecting the optimal viewpoint at a given location. We compare existing approaches in the literature with additional proposed baselines and propose a novel data-driven approach. The result demonstrates the superior performance of our data-driven approach when compared to existing methods, both in controlled simulation experiments and real-world deployment.
@@ I Introduction
As an example of this larger question, this paper delves into the specific scenario of localizing a new agent, such as a ground robot, in a pre-existing map. This application holds considerable practical value, as it obviates the need to re-map an entire building if a suitable map already exists. Furthermore, for seamless collaboration between human-robot teams, the ability to localize mobile devices and robots within a shared map becomes imperative, as highlighted in previous studies [2, 3, 4]. However, accurately localizing a robot in a map created with a head-mounted camera rig introduces its own challenges. These challenges stem from the use of diverse sensor devices and are compounded by significant variations in viewpoint between the mapping trajectory and the operational height of the robot. Such variations result in multiple causes that diminishes the localization performances, such as reduced visual overlap. This issue becomes particularly prominent in the case of ground robots, including quadrupeds, where obstacles such as chairs, tables, and furniture frequently obscure substantial portions of the robot’s environment.
@@
Figure 1: Viewpoint Selection of three methods that run visual localization at the same location with respect to the built map (landmarks in red). The passive strategy of looking forward. and a strategy inspired by [5] to maximize the similarity with the viewing angle towards the landmark during mapping both result in higher localization error than our data-driven viewpoint transformer (VPT) approach.
Many works have studied how to better localize a given image within the map. However, in contrast to always trying to achieve the best from a given viewpoint, robots possess the valuable ability to autonomously select viewpoints. This leads us to investigate whether active viewpoint selection can effectively address the challenges associated with cross-agent visual localization. In the literature, this concept is widely recognized as active perception, and within our specific context, it is referred to as Active Visual Localization. The core objective of Active Visual Localization is to determine the camera pose within an existing map representation of the environment, in order to improve localization accuracy. A common approach involves assessing the localization utility of various viewpoints, typically through metrics like the Fisher Information Metric (FIM), or a combination of hand-crafted heuristics. While extensive work has been directed towards integrating these utility calculations into planning frameworks, a relatively less studied component is how the utility value itself can be better estimated, particularly in scenarios involving the unique challenges discussed earlier.
This work delves into the exploration of an effective utility function to actively enhance visual localization of a robot in a mapped environment. Our primary focus lies in evaluating established viewpoint selection and assessment criteria, while introducing a novel data-driven approach to viewpoint scoring. The key contributions of this paper are:
@@
  • Comparison and thorough evaluation of viewpoint selection methods for the problem of visual localization between heterogeneous agents
  • Real-world validation of our findings by integrating the viewpoint selection into an active viewpoint planner for a robot with an arm-mounted camera
Figure 2: Difference in perspective between a head-mounted sensor rig used for mapping (left) and a ground robot (right) deployed for localization.
@@ II Related Work
Active Vision describes the case where an agent has the ability to move visual sensors with the goal of improving the performance of perceptual tasks [6]. This concept finds application across various domains where visual information is utilized, such as scene exploration [7, 8], data collection [9, 10], inspection [11], and active learning [12, 13]. The case studied in this work focuses on the task of improving visual localization by choosing the most informative viewpoint at a given position, which can be considered a subset of Active Vision, named Active Visual Localization. Numerous existing works have studied how informative metrics can be incorporated into motion planning frameworks with the goal of optimizing robot motion to maximize localization accuracy [14, 15, 16]. Within the scope of this research, our focus is on the specific task of augmenting visual localization by selecting the most informative viewpoint at a given position.
@@
Regarding viewpoint selection in the realm of active vision, early approaches primarily rely on handcrafted metrics to gauge the uncertainty or reliability of a given viewpoint in contributing to the system’s state [17, 18, 19, 20, 21, 22, 23]. Some methods take a different path by deriving evaluation metrics from the metric map and constructing a global utility map. Authors of [24] propose a solution rooted in Fisher information theory [25, 26]. The central task they explore consists of determining the amount of information a viewpoint from a given pose will contribute to the localization process. They develop a novel map representation that enables efficient computation of the Fisher information for 6-DoF visual localization, known as the Fisher Information Field. Similar ideas of using Fisher Information for viewpoint selection have also been explored in recent works [23, 27, 28]. However, it’s important to note that these metrics often rely on heuristics, necessitate the design of specific handcrafted utility functions, and may have limitations in their representation capabilities for diverse and complex scenarios.
@@
Another category of works introduces additional vision tasks to enhance the metric extraction process, notably incorporating semantic information [29]. For instance, the work of [30] aims to improve navigation performance by including semantic information, in order to discern perceptually informative areas of the environment. This kind of work enriches the semantic understanding capability, however highly relies on the performance of the semantic module, and usually requires prior knowledge to link certain semantic classes that clearly correlate to visual informativeness.
@@
A separate line of research focus on using data-driven approaches to active visual localization. These works mostly choose to formulate the problem as a reinforcement learning task, tightly integrating metric determination with robot execution [31, 32]. As an example, [33] trains an information-aware policy to find traversable paths as well as reduce the uncertainty of the environment. The learning-based model significantly enhance the robot’s comprehension of its surroundings. However, it’s worth noting that these models typically demand substantial computational resources for training and may require the simplification of the environment model to prevent over-fitting.
@@ III Method
Figure 3: Overview of the proposed active localization approach The core of our approach is the learning-based viewpoint evaluation model. This model processes input features derived from an established Structure-from-Motion model alongside a camera viewpoint. It predicts the likelihood of the given viewpoint being effective for visual localization. In practice, when deployed, multiple viewpoints are sampled and assessed at a particular 3D location. The viewpoint that receives the highest predicted score is then chosen as the optimal one to execute for the robot.
Problem Statement: Let $\bm{x}=(\bm{p},\bm{q})$, where $\bm{p}\in\mathcal{R}^{3}$ and $\bm{q}\in so(3)$ are the position and orientation of the robot sensor. Given a map representation of the environment $\mathcal{M}$ and a prior of the robot position $\hat{\bm{p}}$, our goal of the active visual localization is to find a viewpoint at that location, i.e., the orientation for the robot sensor, such that the visual localization method returns the most accurate estimation $\bar{\bm{x}}$ at that location, i.e.,
@@
$=\operatorname*{arg\,min}_{\bm{q}}\lVert\textit{loc}(\mathcal{M},\mathcal{O}(\bm{x}))-\bm{x}\rVert$ (2)
Here $\textit{loc}(\cdot)$ refers to different localization methods, which often take the observation $\mathcal{O}$, captured at the current pose $\bm{x}$, and localize it against the given map representation $\mathcal{M}$.
In this paper, we construct $\mathcal{M}$ that combines the landmark point cloud $\mathcal{M}_{l}$ and the Truncated Signed Distance Function (TSDF) $\mathcal{M}_{t}$ of the environment. Any viewpoint selection policy $\uppi(\cdot)$ then takes $\mathcal{M}=(\mathcal{M}_{l},\mathcal{M}_{t})$ as prior knowledge of the environment, and selects a viewpoint for a certain estimated position:
$\bm{q}^{\star}$ $=\uppi(\mathcal{M},\hat{\bm{p}})$ (3)
As an example, one of the baseline methods we implement takes the idea from [24]:
@@
where $v(\cdot)$ is the binary visibility of landmark $i$ and $\textrm{I}_{i}$ is the Fisher Information Metric (FIM) of observing landmark $i$. The visibility $v(\cdot)$ is determined using TSDF map $\mathcal{M}_{t}$.
@@
where $\uppi_{\textrm{max}}$ selects simply the view with the maximum visible landmarks, and $\uppi_{\textrm{angle}}$ uses a stricter visibility criterion inspired by [5] that only considers a landmark visible if its relative location to $\bm{x}$ is similar to those poses from which it was seen during mapping.
We propose a data-driven approach, adhering to a ”sample-and-evaluate” framework as depicted in Figure 3. Our process involves gathering simulated data to train our viewpoint evaluation model, which is then tested through both simulated scenarios and real-world experiments. In the inference phase, we sample multiple viewpoints at a given location. For each viewpoint, we assemble its input feature vector by utilizing the landmark point cloud derived from the Structure-from-Motion (SfM) model. The rest of this section will detail our approaches, with the focus on constructing the feature vectors and the viewpoint evaluation model.
Data-driven viewpoint evaluation For each viewpoint, we collect the following essential information:
  • Its distance to every landmark of the map.
  • The viewing angle between every landmark and the principle axis.
  • The minimum and maximum distance and viewing angle per landmark has, with respect to the camera frame during the mapping stage
@@
  • The number of landmarks in the previously seen angle range
  • Its corresponding DINO [34] appearance features for every landmark
To collect training data, we generate for every viewpoint the above information, as well as the ground truth pose and the result of the visual localization method at that viewpoint. Based on this data, we train our model to classify whether a viewpoint has been localized within an error threshold. In particular, we distinguish between two methods, one is based on Multi-Layer Perceptron (MLP) $\uppi_{MLP}$, and the other one is based on the Transformer $\uppi_{VPT}$. The way of encoding the collected information differs respectively. We design lightweight models in order to maintain online capability. Both kinds of models end with a softmax layer to determine the viewpoint score for classification.
The model based on MLP requires a fixed input dimensionality. However, the number of landmarks varies across different viewpoints, we set a feature aggregation step in the model of $\uppi_{MLP}$. For each information in the list above, we build a histogram to aggregate the corresponding value of all the filtered landmarks. For the pixel coordinates, we aggregate the information in a 2D heatmap instead of a 1D histogram. DINO features cannot be processed in the MLP-based model.
The constraints of fixed input dimensionality also encourages our motivation to investigate the transformer architecture. With the transformer, our inputs can be provided in a per-landmark fashion, which also allows for the inclusion of features that cannot be easily aggregated, such as DINO appearance per landmark.
Upon obtaining the trained model, for a specified point $\bm{p}$, we assess all the sampled viewpoints. Each candidate is allocated a localization score, reflecting its efficacy in visual localization. Subsequently, we select the viewpoint that is awarded the highest score. For instance:
@@
where $\mathbf{F}$ are the input features, and $f_{VPT}(\cdot)$ returns the output score from the viewpoint transformer (VPT) model.
@@ IV Experiments
We develop a comprehensive pipeline for data collection aimed at training, validation, and testing across simulated and real-world settings, and we carry out a variety of experiments. Additionally, we implement several baseline methods for comparison. The data collection and model training is mostly done within the simulation, details can be found in IV-A and 4. We present all baseline methods alongside our evaluation strategy in IV-C. Finally, our real-world experiment findings are detailed in IV-D.
@@ IV-A Data Generation
Simulated Scenario We focus on indoor scenarios, and choose a selection of scenes from the Habitat-Matterport 3D (HM3D) dataset [35]. HM3D provides high-quality 3D reconstructions of real-world indoor environments with textured 3D meshes. We selected nine scenes, as depicted in Figure 4, and imported them into NVIDIA Isaac Sim. Subsequently, we manually captured controlled trajectories using a simulated camera setup that mimics the Microsoft HoloLens 2 [36], ensuring the camera was positioned at a height akin to that of a human.
Data Collection To acquire $\mathcal{M}_{t}$, we utilize the depth images captured by sensors and conduct TSDF (Truncated Signed Distance Field) integration [^1] Isaac Sim offers a direct method to ascertain the occupancy details of a 3D scene through its built-in functionality, which we also employ as an alternative approach.. Both the generated meshes and the occupancy maps can be used for occlusion and collision checking.
To obtain $\mathcal{M}_{l}$, we use mapping and localization frameworks from hloc and COLMAP [37, 38] to extract 2D local features from images and build a 3D landmark point cloud. The collected images from the simulated HoloLens 2 are fed into the pipeline to build the 3D landmarks map. The per-landmark feature vector $\mathbf{F}$ are also created and attached to each landmark at this stage.
For running visual localization, We mainly use the localization module from hloc as the framework for $\textit{loc}(\cdot)$ in 2.
To recover the scale of the map, one option would be to create the reconstruction with known camera poses. However, this would not be representative of how this process could be done using real hardware, where poses would be estimated, for example, from a visual SLAM module. In our experiment, all maps are created without the exact camera poses. Instead, we build the reconstructions from the images alone and then using RANSAC to estimate the scale with respect to the ground-truth poses.
To collect training data point in the simulator, we create random camera paths at a height characteristic of robots, using a single RGB-D camera to mimic an onboard robot sensor. At each waypoint, we capture a set of viewpoint samples and store the images along with their exact viewpoint poses as ground truth. As illustrated in 1 and 2, we then employ the localization module for each sample, calculating the discrepancy between the estimated pose and the actual ground truth pose.
@@
TABLE I: Evaluation of viewpoint strategies in the simulated testing environments using both SuperPoint(Left) and SIFT(Right) features. Recall percentages are shown at varying distance and orientation thresholds, highlighting the best and second-best performing methods. Rows marked as ‘with occlusion filter’ are those where landmarks are filtered with the occlusion. The bottom row is an oracle method that selects the viewpoint with smallest possible error among all samples.